OpenStudy (anonymous):

factor trinomial 14y^2-12y+3

OpenStudy (mathstudent55):

Again, the first step is to look for a common factor. There isn't one.

OpenStudy (mathstudent55):

Second step, multiply ac. What do you get?

OpenStudy (anonymous):

why isn;t there one?

OpenStudy (mathstudent55):

14, -12, and 3 are not all divisible by any number other than 1.

OpenStudy (anonymous):

ok

OpenStudy (mathstudent55):

So there is no common factor that can be factored out of all three terms.

OpenStudy (mathstudent55):

Now we move to the next step. Multiply ac

OpenStudy (anonymous):

42

OpenStudy (mathstudent55):

Good. Now what two factors of 42 add to -12?

OpenStudy (mathstudent55):

Since 42 is positive, and -12 is negative, both factors have to be negative. What two numbers multiply to +42 and add to -12?

OpenStudy (compassionate):

It's not factorable using real numbers. 14y^2-12y+3 2y(7y - 6) + 3

OpenStudy (anonymous):

cant figure it out

OpenStudy (mathstudent55):

-42 * (-1) = 42, but -42 + (-1) = -43, not -12 -21 * (-2) = 42, but -21 + (-2) = -23, not -12 -14 * (-3) = 42, but -14 + (-3) = -17, not -12 -7 * (-6) = 42, but -7 + (-6) = -13, not -12 There are no two factors of 42 that add up to -12, so this trinomial is not factorable.

OpenStudy (anonymous):

i knew it!! thanks